OpenStudy (anonymous):

the angle measures of a triangle are x, 2x , and 12x degrees. find each angle measure.

OpenStudy (anonymous):

The sum in any triangle is 180 degrees. x + 2x + 12x = 180 15x = 180 x = 180/15 x=12 angle 1 = 12 deg angle 2 = 2*12 = 24 deg angle 3 = 12*12 = 144 deg